OpenStudy (anonymous):

Solve x – 2y = 5 using the substitution method. 2x – 4y = 1

OpenStudy (mrnood):

x – 2y = 5 first re-arrange this equation so it is x=???

OpenStudy (anonymous):

yes both the equations are cancelling

OpenStudy (mrnood):

Hmmm These equations are equivalent to y=(x-5)/2 and y=(2x-1)/4 both these lines have a slope of 1/2 and are therefore parallel Draw your own conclusions from that...